Given a population of 5000 scores with mean= 86 and SD= 10. How many scores are above 116

1
Expert's answer
2022-04-28T12:05:16-0400
"P(X>116)=1-P(Z\le\dfrac{116-86}{10})"

"=1-P(Z\le3)\approx0.001350"

"0.001350(5000)=7"

"7" scores